# On the long-time limit of the mean curvature flow in closed manifolds

### v1

@@ 2.2. Ilmanen’s regularity theorem
By Lemma 2.5, item (1) holds because this homology group is closed under limits. Now for the other parts we first note that because the convergence to the $\Sigma_{j}$ is multisheeted away from a discrete set of points, each sheet in the convergence corresponds to a connected component of the boundary of the tubular neighborhood of the limit (or, in other words, a connected component of its normal bundle). If $\Sigma_{j}$ is $1$-sided there is only one connected component, but about a point locally there are two components giving item (2). For item (3) by genus of nonoriented $\Sigma_{j}$ we refer to the demigenus and, since this may be somewhat less familiar, first we consider the case all the $\Sigma_{j}$ are orientable. In this case because all of the limit surfaces $\Sigma_{j}$ are orientable each sheet over $\Sigma_{j}$ in the convergence is homeomorphic to $\Sigma_{j}$ away from a finite set of points following the discussion above; then the claim follows easily from Lemma 2.6. In case some are nonorientable we note by the classification of surfaces they must be homeomorphic to a connected sum of projective planes, say $k_{j}$ of them. The demigenus of such a connect sum is $k_{j}$ and the double cover of such a connect sum, corresponding to a single sheet in the convergence, happens to have genus $k_{j}-1$. Because these numbers agree we obtain the formula in item (3). ∎
@@
From the previous section we see that if the limit of an almost regular flow is nonempty it converges to a collection of smooth minimal surfaces, possibly with multiplicity. We also obtain some rough topological information about the limit. Even so it is still often desirable to specifically produce stable minimal surfaces because they carry more structure; for instance as a basic but important example we recall that stable surfaces in PSC $3$–manifolds must be diffeomorphic to $S^{2}$ or ${\mathbb{R}}P^{2}$ but just in $S^{3}$ there are minimal surfaces of arbitrarily large genus. In this section, we show that with appropriately timed small isotopies, we can construct piecewise flows whose long-time limits are either empty or stable minimal surfaces.
We consider first the Ricci positive case. As promised in the introduction, arguing in this case turns out to be simple because in this case the avoidance principle says two disjoint flows actually "repel" In fact we don’t even need the regularity statements from the previous statement to proceed and indeed our argument works in all dimensions.
@@
Where $M$ and $N$ are as in the statement of the theorem, suppose that the flow $M_{t}$ of $M$ doesn’t go extinct in finite time. Since the flow is gradient of the area functional, we may pick a sequence of times $t_{i}\to\infty$ for which $\int_{M_{t_{i}}}H^{2}\to 0$. After potentially passing to a subsequence the sequence $M_{t_{i}}$ of surfaces limits to a varifold $\Sigma$ which by the choice of $t_{i}$ must be stationary. Now consider a one sided perturbation $M^{\prime}$ of $M$, for which say $d(M,M^{\prime})>\delta$ for some $\delta>0$. Since $N$ is compact and Ricci positive, there exists some $\Lambda>0$ for which the Ricci curvature of $N$ is bounded below by $\Lambda$. Hence by the avoidance principle, Theorem 2.1 above, we have that the distance between $M^{\prime}_{t}$ and $\Sigma$ is bounded below by $\delta e^{\Lambda t}$ for all times; because $N$ is compact so has bounded diameter for $t$ sufficiently large this gives that the level set flow $M^{\prime}_{t}$ of $M^{\prime}$ must be empty. ∎
We point out that as a consequence of the above, that in an orientable $3$–manifold every $2$–cycle can be represented by a properly embedded $2$–sided surface (see, for instance, Lemma 3.6 in [22]), and Proposition 3.7 we have another proof of the following classical fact:
@@
$H_{2}(N,{\mathbb{R}})$ of a compact orientable $3$–manifold $N$ which admits a metric of positive Ricci curvature vanishes.
Indeed, by Hodge theory and the Bochner formula $H^{1}(N)$ vanishes so the claim follows by Poincare duality. Now in the case $N$ doesn’t have Ricci positive curvature the avoidance principle doesn’t give as strong information and indeed it is easy to concoct examples where two disjoint flows may converge to each other as $t\to\infty$. For instance let $N$ be a compact manifold for which $H_{2}(N,{\mathbb{Z}})$ is nontrivial (so cannot have positive Ricci curvature by the above). By the direct method in a nontrivial class of $H_{2}(N,{\mathbb{Z}})$ there is an area minimizing surface $\Sigma$, and by perturbing the metric to be bumpy we may even suppose it is strictly area minimizing. By considering disjoint nearby surfaces (say graphical over $\Sigma$) their flows must converge to $\Sigma$ in the long term.
With this in mind we now move on to Theorem 1.4. Recall the definition of piecewise generic flow, Definition 2.7, and note that after potentially perturbing $M$ be a small amount without loss of generality we may suppose we are in the setting of Corollary 1.1. If $M_{t}$ goes extinct or all the minimal surfaces $\Sigma_{j}$ it subconverges to, defined as in the previous section, are stable minimal surfaces, then we are done. Otherwise, we apply isotopes to $M_{t}$ at a smooth time far along the flow which ensures the perturbation must not converge to the same surfaces after restarting the flow:
@@
Let $M$ be a $2$–sided properly embedded surface in a closed compact Riemannian $3$–manifold $(N^{3},g)$, and suppose $M_{t}$ is an almost regular flow out of $M$ which can be approximated by generic flows (if not generic itself, the inner or outermost flow from $M$). Suppose $M_{t_{i}}$ converges to $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as varifolds, and one of $\Sigma_{j}$ is unstable. Then for sufficiently large $t_{i}$ we may isotope $M_{t_{i}}$ so that any almost regular flow starting from the perturbed surface can not enter the tubular neighborhood of that unstable $\Sigma_{j}$.
@@
Suppose $\Sigma_{1}$ is an unstable minimal surface, and in the following, we simply call it $\Sigma$. We first consider the situation that $\Sigma$ is unstable and $2$-sided. Then we can find a unit normal vector field $\nu$ over it, and we may assume $\varphi$ is the first eigenfunction of the linearized operator $L_{\Sigma}$. Let us use $(x,s)\in\Sigma\times(-\tau,\tau)$ to parametrize a tubular neighborhood $\mathcal{N}_{\tau}$ of $\Sigma$ by $(x,s)\to\operatorname{exp}_{x}(s\varphi(x)\nu(x))$. We first define an isotope $\Psi_{i}$ of the ambient manifold as follows: we define the vector field $\partial_{s}$ in $\mathcal{N}_{\tau/2}$ using the coordinate $(x,s)$, then we smoothly extend it to 00 outside $\mathcal{N}_{\tau}$ to get a smooth vector field $W$. Then we define $\{\Psi(w)\}_{w\in{\mathbb{R}}}$ to be the isotopy generated by the vector field $W$.
@@
Let us choose a very small $\delta>0$, such that $\mathcal{N}_{\delta}(\Sigma)$ is completely contained in the domain given by the parametrization $(x,s)\to\operatorname{exp}_{x}(s\varphi(x)\nu(x))$ for $s\in(-\tau/10,\tau/10)$, and $\Psi(\tau/2)\mathcal{N}_{\delta}(\Sigma)$ will completely leave $\mathcal{N}_{\delta}(\Sigma)$. Following Lemma 3.5, we may assume $M_{t_{i}}\subset\bigcup_{j=1}^{k}\mathcal{N}_{\delta}(\Sigma_{j})$ for sufficiently large $t_{i}$. Then $\tilde{M}_{t_{i}}:=\Psi(\tau/2)M_{t_{i}}$ must be completely disjoint from $\mathcal{N}_{\delta}(\Sigma)$, and this is a $2$–sided flow from which we may safely restart the flow from. Recalling that the support of Brakke flows are weak set flows, it’s easy to see that any Brakke (and hence almost regular) flow starting from such $M^{\prime}:=\tilde{M}_{t_{i}}$ will never enter $\mathcal{N}_{\delta}(\Sigma)$ again by the avoidance principle. To see this, note that the domain generated by the parametrization $(x,s)\to\operatorname{exp}_{x}(s\varphi(x)\nu(x))$ for $s\in(-\tau/10,\tau/10)$ has compact, smooth and outwardly mean convex boundary, with two components, which we denote by $\Sigma_{\pm\tau/10}$. By the strict mean convexity and the compactness there exists $\delta,\eta>0$ so that the smooth mean curvature flow of the boundary exists on $[0,\delta)$ and the distance between $\Sigma_{\pm\tau/10}$ and $(\Sigma^{\prime}_{\pm\tau/10})_{\delta}$ is bounded below by $\eta$. The avoidance principle says $(\Sigma_{\pm\tau/10})_{t}$ and $M^{\prime}_{t}$ are disjoint on $[0,\delta)$, and because $\eta>0$ we can iterate this argument over the intervals $[\delta,2\delta)$ et. cetera to get the claim.
@@
Finally, we briefly describe the situation that $\Sigma$ is unstable and $1$–sided. The argument is similar but now the tubular neighborhood of $\Sigma$ is only parametrized by $\overline{\Sigma}\times(0,\tau)\bigcup\Sigma$, where $\overline{\Sigma}$ is the oriented double cover of $\Sigma$. In this case, we can still construct the isotopy to push the flow outside this tubular neighborhood, and verbatim arguments validate all the previous arguments. ∎
@@
We may clearly restart the flow after the perturbation in Lemma 4.2, but apriori the set of times for which we perturb may accumulate to a finite time; we claim that only finitely many perturbations are required though to ensure that afterwards none of the $\Sigma_{j}$ could be unstable. Suppose $\Sigma$ is an unstable minimal surface that the flow was perturbed to avoid by the lemma; then if the flow $M_{t}$ starting from this perturbation converges to another unstable minimal surface it must converge to one which is disjoint from $\Sigma$ and hence disjoint from the component of the flow which was just perturbed – but by the avoidance principle argument above the flow of this connected component may never converge to any such unstable minimal surface.
Now because we can take the parameter $\delta$ in the proof above as small as we wish we may suppose the area of the piecewise mean curvature flow is bounded by, say, twice the area of $M$ after the $k$-th perturbation/iteration of Lemma 4.2 for any $k\in\mathbb{N}$. In particular by the clearing out lemma only finitely many components of $M_{t}$ at a given smooth time might go on to flow to a minimal surface, with the rest must quickly being exhausted; in fact by this we may crudely uniformly upper bound (independent of time) the number of such components. Applying the reasoning above to each such component then gives what we want.∎
@@
In this section (aside from the last application, where we use Theorem 1.1), we consider various circumstances where we can show that the piecewise almost regular flows from the last section will be nonempty as $t\to\infty$ and say something about the topology of the limit. The first is on the existence of stable minimal spheres when $\pi_{2}(N)\neq 0$.
@@
First we note that by the sphere theorem in $3$–manifold topology [40] we can find, using $N$ is orientable, a homotopically nontrivial embedded $2$–sphere $M$. Of course $M$ is itself orientable, and we may use it as initial data in Theorem 1.4. By considering the approximating surgery flows $M_{t}$ during smooth times will consist of a collection (possibly multiple, if there are nontrivial neckpinches) of embedded $2$–spheres. Now, either the flow from the conclusion will terminate in finite time or converge to a stable minimal surface $\Sigma$ which, by the nature of the convergence discussed in Theorem 1.1 and the proof of Proposition 3.7, must be a collection of embedded stable minimal spheres or projective planes because each component for large (smooth) times is covered $k\geq 1$ times by $M_{t}$ for large $t$, away from mentioned bridge points and each component of this is an embedded $2$–sphere. Picking any of these components then gives the statement.
So, it suffices to rule out the flow going extinct in finite time. Supposing this is actually the case, considering the approximating surgery flow one can show that $M$ is isotopic to a "marble" tree as in [3] or similarly [37], taking some care because in our setting the flow may not be mean convex even if the high curvature regions are. Because in the mean curvature flow with surgery for large parameters the high curvature regions all bound handlebodies one can show such marble trees bound $3$–balls in $N$ giving that $M$ is in fact homotopically trivial, leading to a contradiction.∎
We remind the reader the next corollary is on the existence of stable minimal tori in the aspherical case. Corollary 1.7 in fact follows from Corollary 1.8: it isn’t hard to see by Dehn’s lemma and the trivial $\pi_{2}$ assumption that if the torus $T$ isn’t algebraically incompressible then it bounds a 3 cycle. That being said we give the proof below for the sake of exposition, because it is somewhat simpler than the more general statement.
@@
Denote by $M$ the torus in the statement. Because the torus is orientable and $N$ is assumed oriented we have that $M$ is $2$–sided, so that we may apply Theorem 1.4 to it. First we argue that the flow must not go extinct. Supposing it does, we note that approximating surgery flows go extinct as well. Considering one of these and denoting it by $S_{t}$, we note that high curvature regions which are discarded certainly bound $3$–cycles, homeomorphic to either $3$–balls or solid tori $\simeq S^{1}\times D^{2}$, since different time slices along a smooth mean curvature flow are homologous this gives that $M$ is homologous to $S_{t}$ for all $t>0$ and hence bounds a $3$–cycle, contradicting that $M$ is assumed to be homologically nontrivial.
@@
Next, we note that for any finite smooth time $t$, $M_{t}$ cannot be given by a union of embedded spheres $P_{1},\ldots,P_{k}$. To see this, because $\pi_{2}(N)$ is trivial each of the $P_{i}$ are nullhomotopic and hence are the boundary of a $3$–cycle, so since $M_{t}$ is homologous to $M$ this gives that $M$ is homologically trivial giving a contradiction. In particular, a connected component of $M_{t}$ at a smooth time must be homeomorphic to a torus. Now, because the flow doesn’t go extinct as before we may produce a sequence of times $t_{i}\to\infty$ for which $M_{t_{i}}$ converge to an embedded stable minimal surface $\Sigma$, and we claim one of the connected components of $\Sigma$ is a torus or a Klein bottle (which is the nonorientable surface covered by the torus) – note this doesn’t follow immediately from the case that a component of $M_{t}$ is a torus. For example, one can take a standardly embedded torus in ${\mathbb{R}}^{3}$, tighten the "donut hole" and then push the sides of the inner cylinder along the outer sides of the donut so in this fashion covering the sphere twice, away from bridge points. Of course if one of the components is a torus or Klein bottle we are done, so it suffices to rule out the case none of them are. By Proposition 3.7 in this case all the components of $\Sigma$ are spheres or projective planes.
@@
Suppose first all the components are spheres: then by assumption they are nullhomotopic and so each of these spheres then could be homotoped to be within coordinate patches homeomorphic to open balls. As these patches are homeomorphic to an open ball in ${\mathbb{R}}^{3}$ all surfaces contained within are nullhomologous. Using that homotopic surfaces are homologous as before this implies $M$ itself is nullhomologous, giving a contradiction. If some of the components are projective planes we can argue similarly, because the boundary of their tubular neighborhoods are diffeomorphic to spheres. ∎
@@
As the reader recalls from the introduction, the next statement concerns the existence of stable minimal surfaces in the presence of algebraically incompressible initial data, which is to say the inclusion map on fundamental groups injects; by Dehn’s lemma, algebraically incompressible surfaces are incompressible in the standard sense i.e. that there are no compressing discs in $N$ along $M$. This topological assumption is arguably most naturally suited to how the topology of a flow may change across singularities, at least generic ones. This is because the effect of flowing through a neckpinch is essentially to glue a thickened disc along the flow of the corresponding domain, and so if the surface is incompressible only spheres can be pinched off.
@@
Supposing that $M=\iota(\Sigma)$ is as in the statement, we consider as before the approximating surgery flow $S_{t}$ to $M_{t}$. We first claim that at any smooth time a given connected component of it is either homeomorphic to either a sphere or $M$. Supposing not and let $T$ be the smallest time, necessarily right after a surgery, that this isn’t the case – for a fixed choice of approximating surgery flow this time is well defined. Then it’s easy to see that the cross section curve $\gamma$ at one of the necks where surgery was performed on $S_{T}^{-}$ (the presurgery surface) is homotopically nontrivial in $M_{T}$. The surgery shows that $\gamma$ bounds a disc in $N$ though, so that $S_{T}^{-}$ is not algebraically incompressible. Since $T$ was the least such time this implies that $M$ itself is not algebraically incompressible, which gives a contradiction.

### v2

@@ 2.2. Ilmanen’s regularity theorem
By Lemma 2.5, item (1) holds because this homology is preserved under limits. Now for the other parts, we first note that because the convergence to the $\Sigma_{j}$ is multisheeted away from a discrete set of points, each sheet in the convergence corresponds to a connected component of the boundary of the tubular neighborhood of the limit (or, in other words, a connected component of its unit normal bundle which, for a surface $\Sigma$, in this case will have either one or two connected components depending on the orientability of $\Sigma$). If $\Sigma_{j}$ is $1$-sided, there is only one connected component, but about a point locally there are two components, giving item (2). For item (3), by genus of nonoriented $\Sigma_{j}$ we refer to the demigenus and, since this may be somewhat less familiar, first we consider the case where all the $\Sigma_{j}$ are orientable. In this case, because all of the limit surfaces $\Sigma_{j}$ are orientable, each sheet over $\Sigma_{j}$ in the convergence is homeomorphic to $\Sigma_{j}$ away from a finite set of points, following the discussion above; then the claim follows easily from Lemma 2.6. In case some are nonorientable, we note by the classification of surfaces that they must be homeomorphic to a connected sum of projective planes, say $k_{j}$ of them. The demigenus of such a connect sum is $k_{j}$ and the double cover of such a connect sum, corresponding to a single sheet in the convergence, happens to have genus $k_{j}-1$. Because these numbers agree, we obtain the formula in item (3). ∎
@@
From the previous section, we see that if the limit of an almost regular flow is nonempty, it converges to a collection of smooth minimal surfaces, possibly with multiplicity. We also obtain some rough topological information about the limit. Even so, it is still often desirable to specifically produce stable minimal surfaces because they carry more structure; for instance, as a basic but important example, we recall that stable surfaces in PSC 33–manifolds must be diffeomorphic to $S^{2}$ or ${\mathbb{R}}P^{2}$, but just in $S^{3}$ there are minimal surfaces of arbitrarily large genus. In this section, we show that with appropriately timed small isotopies, we can construct piecewise flows whose long-time limits are either empty or stable minimal surfaces.
We consider first the Ricci positive case. As promised in the introduction, arguing in this case turns out to be simple because, in this case, the avoidance principle says two disjoint flows actually "repel." In fact, we don’t even need the regularity statements from the previous statement to proceed, and indeed, our argument works in all dimensions.
@@
Where $M$ and $N$ are as in the statement of the theorem, suppose that the flow $M_{t}$ of $M$ doesn’t go extinct in finite time. Since the flow is gradient of the area functional, we may pick a sequence of times $t_{i}\to\infty$ for which $\int_{M_{t_{i}}}H^{2}\to 0$. After potentially passing to a subsequence, the sequence $M_{t_{i}}$ of surfaces limits to a varifold $\Sigma$ which, by the choice of $t_{i}$, must be stationary. Now consider a one sided perturbation $M^{\prime}$ of $M$, for which say $d(M,M^{\prime})>\delta$ for some $\delta>0$. Since $N$ is compact and Ricci positive, there exists some $\Lambda>0$ for which the Ricci curvature of $N$ is bounded below by $\Lambda$. Hence by the avoidance principle, Theorem 2.1 above, we have that the distance between $M^{\prime}_{t}$ and $\Sigma$ is bounded below by $\delta e^{\Lambda t}$ for all times; because $N$ is compact so has bounded diameter for $t$ sufficiently large this gives that the level set flow $M^{\prime}_{t}$ of $M^{\prime}$ must be empty. ∎
We point out that, as a consequence of the above, in an orientable 33–manifold every 22–cycle can be represented by a properly embedded 22–sided surface (see, for instance, Lemma 3.6 in [Hatcher07_Notes3Mfd]), and Proposition 3.7 we have another proof of the following classical fact:
@@
$H_{2}(N,{\mathbb{R}})$ of a compact orientable 33–manifold NN which admits a metric of positive Ricci curvature vanishes.
Indeed, by Hodge theory and the Bochner formula, $H^{1}(N)$ vanishes, so the claim follows by Poincaré duality. Now, in the case $N$ doesn’t have Ricci positive curvature, the avoidance principle doesn’t give as strong information, and indeed it is easy to concoct examples where two disjoint flows may converge to each other as $t\to\infty$. For instance, let $N$ be a compact manifold for which $H_{2}(N,{\mathbb{Z}})$ is nontrivial (so cannot have positive Ricci curvature by the above). By the direct method in a nontrivial class of $H_{2}(N,{\mathbb{Z}})$, there is an area minimizing surface $\Sigma$, and by perturbing the metric to be bumpy, we may even suppose it is strictly area minimizing. By considering disjoint nearby surfaces (sa,y graphical over $\Sigma$), their flows must converge to $\Sigma$ in the long term.
With this in mind, we now move on to Theorem 1.4. Recall the definition of piecewise generic flow, Definition 2.7, and Theorem 1.1. If $M_{t}$ goes extinct or all the minimal surfaces $\Sigma_{j}$ it subconverges to, defined as in the previous section, are stable minimal surfaces, then we are done. Otherwise, we apply isotopes to $M_{t}$ at a smooth time far along the flow, which ensures the perturbation must not converge to the same surfaces after restarting the flow:
@@
Let $M$ be a 22–sided properly embedded surface in a closed compact Riemannian 33–manifold $(N^{3},g)$, and suppose $M_{t}$ is an almost regular flow out of $M$ which can be approximated by generic flows (if not generic itself, the inner or outermost flow from $M$). Suppose $M_{t_{i}}$ converges to $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as varifolds, and one of $\Sigma_{j}$ is unstable. Then, for sufficiently large $t_{i}$, we may isotope $M_{t_{i}}$ so that any almost regular flow starting from the perturbed surface can not enter the tubular neighborhood of that unstable $\Sigma_{j}$.
@@
Suppose $\Sigma_{1}$ is an unstable minimal surface, and in the following, we simply call it $\Sigma$. We first consider the situation that $\Sigma$ is unstable and $2$-sided. Then we can find a unit normal vector field $\nu$ over it, and we may assume $\varphi$ is the first eigenfunction of the linearized operator $L_{\Sigma}$. Let us use $(x,s)\in\Sigma\times(-\tau,\tau)$ to parametrize a tubular neighborhood $\mathcal{N}_{\tau}$ of $\Sigma$ by $(x,s)\to\operatorname{exp}_{x}(s\varphi(x)\nu(x))$. We first define an isotope $\Psi_{i}$ of the ambient manifold as follows: we define the vector field $\partial_{s}$ in $\mathcal{N}_{\tau/2}$ using the coordinate $(x,s)$, then we smoothly extend it to 0 outside $\mathcal{N}_{\tau}$ to get a smooth vector field $W$. Then we define $\{\Psi(w)\}_{w\in{\mathbb{R}}}$ to be the isotopy generated by the vector field $W$.
@@
Let us choose a very small $\delta>0$, such that $\mathcal{N}_{\delta}(\Sigma)$ is completely contained in the domain given by the parametrization $(x,s)\to\operatorname{exp}_{x}(s\varphi(x)\nu(x))$ for $s\in(-\tau/10,\tau/10)$, and $\Psi(\tau/2)\mathcal{N}_{\delta}(\Sigma)$ will completely leave $\mathcal{N}_{\delta}(\Sigma)$. Following Lemma 3.5, we may assume $M_{t_{i}}\subset\bigcup_{j=1}^{k}\mathcal{N}_{\delta}(\Sigma_{j})$ for sufficiently large $t_{i}$. Then $\tilde{M}_{t_{i}}:=\Psi(\tau/2)M_{t_{i}}$ must be completely disjoint from $\mathcal{N}_{\delta}(\Sigma)$, and this is a 22–sided flow from which we may safely restart the flow from. Recalling that the supports of Brakke flows are weak set flows. It’s easy to see that any Brakke (and hence almost regular) flow starting from such $M^{\prime}:=\tilde{M}_{t_{i}}$ will never enter $\mathcal{N}_{\delta}(\Sigma)$ again, by the avoidance principle. To see this, note that the domain generated by the parametrization $(x,s)\to\operatorname{exp}_{x}(s\varphi(x)\nu(x))$ for $s\in(-\tau/10,\tau/10)$ has compact, smooth and outwardly mean convex boundary, with two components, which we denote by $\Sigma_{\pm\tau/10}$. By the strict mean convexity and the compactness there exists $\delta,\eta>0$ so that the smooth mean curvature flow of the boundary exists on $0,\delta)$ and the distance between $\Sigma_{\pm\tau/10}$ and $(\Sigma^{\prime}_{\pm\tau/10})_{\delta}$ is bounded below by $\eta$. The avoidance principle says $(\Sigma_{\pm\tau/10})_{t}$ and $M^{\prime}_{t}$ are disjoint on $0,\delta)$, and because $\eta>0$ we can iterate this argument over the intervals $\delta,2\delta)$ et. cetera to get the claim.
@@
Finally, we briefly describe the situation that $\Sigma$ is unstable and 11–sided. The argument is similar, but now the tubular neighborhood of $\Sigma$ is only parametrized by $\overline{\Sigma}\times(0,\tau)\bigcup\Sigma$, where $\overline{\Sigma}$ is the oriented double cover of $\Sigma$. In this case, we can still construct the isotopy to push the flow outside this tubular neighborhood, and verbatim arguments validate all the previous arguments. ∎
@@
We may clearly restart the flow after the perturbation in Lemma 4.2, but a priori the set of times for which we perturb may accumulate to a finite time; we claim that only finitely many perturbations are required, though, to ensure that afterwards none of the $\Sigma_{j}$ could be unstable. Suppose $\Sigma$ is an unstable minimal surface that the flow was perturbed to avoid by the lemma; then if the flow $M_{t}$ starting from this perturbation converges to another unstable minimal surface it must converge to one which is disjoint from $\Sigma$ and hence disjoint from the component of the flow which was just perturbed – but by the avoidance principle argument above the flow of this connected component may never converge to any such unstable minimal surface.
Recall that the clearing out lemma [Ilmanen94_EllipReg, 12.2] asserts that if the area of a flow $M_{t_{0}}$ inside $B_{R}(x)$ is sufficiently small, then after a short time, $M_{t_{0}+cR^{2}}$ will be disjoint from $B_{R/2}(x)$. Now, because we can take the parameter $\delta$ in the proof above as small as we wish, we may suppose the area of the piecewise mean curvature flow is bounded by, say, twice the area of $M$ after the $k$-th perturbation/iteration of Lemma 4.2 for any $k\in\mathbb{N}$. In particular, by the clearing out lemma, only finitely many components of $M_{t}$ at a given smooth time might go on to flow to a minimal surface, and the rest of those with small area must quickly be exhausted. By this, we may crudely uniformly upper bound (independent of time) the number of such components. Alternatively, one may show there is a lower area bound on minimal surfaces of $N$ with bounded genus using [White18_CptMinSurf3-Mfd] and use this. Applying the reasoning above to each such component then gives what we want.∎
@@
In this section (aside from the last application, where we use Theorem 1.1), we consider various circumstances where we can show that the piecewise almost regular flows from the last section will be nonempty as $t\to\infty$ and say something about the topology of the limit. The first is on the existence of stable minimal spheres/projective planes when $\pi_{2}(N)\neq 0$.
@@
First we note that by the sphere theorem in 33–manifold topology [Papakyriakopoulos57_Dehn] we can find, using $N$ is orientable, a homotopically nontrivial embedded 22–sphere $M$. Of course, $M$ is itself orientable, and we may use it as initial data in Theorem 1.4. By considering the approximating surgery flows $M_{t}$ during smooth times will consist of a collection (possibly multiple, if there are nontrivial neckpinches) of embedded 22–spheres. Now, either the flow from the conclusion will terminate in finite time or converge to a stable minimal surface $\Sigma$ which, by the nature of the convergence discussed in Theorem 1.1 and the proof of Proposition 3.7, must be a collection of embedded stable minimal spheres or projective planes because each component for large (smooth) times is covered $k\geq 1$ times by $M_{t}$ for large $t$, away from mentioned bridge points and each component of this is an embedded 22–sphere. Picking any of these components then gives the statement.
So, it suffices to rule out the flow going extinct in finite time. Supposing this is actually the case, considering the approximating surgery flow, one can show that $M$ is isotopic to a "marble" tree as in [BuzanoHaslhoferHershkovits21_Moduli] or similarly [Mramor18_Finiteness], taking some care because in our setting the flow may not be mean convex even if the high curvature regions are. Because in the mean curvature flow with surgery for large parameters, the high curvature regions all bound handlebodies, one can show such marble trees bound 33–balls in $N$, giving that $M$ is in fact homotopically trivial, leading to a contradiction.∎
We remind the reader that the next corollary is on the existence of stable minimal tori in the aspherical case. Corollary 1.7 in fact follows from Corollary 1.8: it isn’t hard to see by Dehn’s lemma and the trivial $\pi_{2}$ assumption that if the torus $T$ isn’t algebraically incompressible, then it bounds a 33–cycle. That being said, we give the proof below for the sake of exposition, because it is somewhat simpler than the more general statement.
@@
Denote by $M$ the torus in the statement. Because the torus is orientable and $N$ is assumed oriented, we have that $M$ is 22–sided, so that we may apply Theorem 1.4 to it. First, we argue that the flow must not go extinct. Supposing it does, we note that approximating surgery flows go extinct as well. Considering one of these and denoting it by $S_{t}$, we note that high curvature regions which are discarded certainly bound 33–cycles, homeomorphic to either 33–balls or solid tori $\simeq S^{1}\times D^{2}$, since different time slices along a smooth mean curvature flow are homologous this gives that $M$ is homologous to $S_{t}$ for all $t>0$ and hence bounds a 33–cycle, contradicting that $M$ is assumed to be homologically nontrivial.
@@
Next, we note that for any finite smooth time $t$, $M_{t}$ cannot be given by a union of embedded spheres $P_{1},\ldots,P_{k}$. To see this, because $\pi_{2}(N)$ is trivial, each of the $P_{i}$ is nullhomotopic, and hence is the boundary of a 33–cycle, so since $M_{t}$ is homologous to $M$, this gives that $M$ is homologically trivial, giving a contradiction. In particular, a connected component of $M_{t}$ at a smooth time must be homeomorphic to a torus. Now, because the flow doesn’t go extinct as before we may produce a sequence of times $t_{i}\to\infty$ for which $M_{t_{i}}$ converge to an embedded stable minimal surface $\Sigma$, and we claim one of the connected components of $\Sigma$ is a torus or a Klein bottle (which is the nonorientable surface covered by the torus) – note this doesn’t follow immediately from the case that a component of $M_{t}$ is a torus. For example, one can take a standardly embedded torus in ${\mathbb{R}}^{3}$, tighten the "donut hole" and then push the sides of the inner cylinder along the outer sides of the donut so in this fashion covering the sphere twice, away from bridge points. Of course, if one of the components is a torus or Klein bottle, we are done, so it suffices to rule out the case that none of them are. By Proposition 3.7, in this case, all the components of $\Sigma$ are spheres or projective planes.
@@
Suppose first all the components are spheres: then by assumption they are nullhomotopic, and so each of these spheres could then be homotoped to be within coordinate patches homeomorphic to open balls. As these patches are homeomorphic to an open ball in ${\mathbb{R}}^{3}$, all surfaces contained within are nullhomologous. Using that homotopic surfaces are homologous as before, this implies $M$ itself is nullhomologous, giving a contradiction. If some of the components are projective planes, we can argue similarly, because the boundary of their tubular neighborhoods is diffeomorphic to spheres. ∎
@@
As the reader recalls from the introduction, the next statement concerns the existence of stable minimal surfaces in the presence of algebraically incompressible initial data, which is to say the inclusion map on fundamental groups injects; by Dehn’s lemma, algebraically incompressible surfaces are incompressible in the standard sense, i.e., that there are no compressing discs in $N$ along $M$. This topological assumption is arguably most naturally suited to how the topology of a flow may change across singularities, at least generic ones. This is because the effect of flowing through a neckpinch is essentially to glue a thickened disc along the flow of the corresponding domain, and so if the surface is incompressible, only spheres can be pinched off.
@@
Supposing that $M=\iota(\Sigma)$ is as in the statement, we consider, as before, the approximating surgery flow $S_{t}$ to $M_{t}$. We first claim that at any smooth time, a given connected component of it is either homeomorphic to either a sphere or $M$. Supposing not, and let $T$ be the smallest time, necessarily right after a surgery, that this isn’t the case – for a fixed choice of approximating surgery flow, this time is well defined. Then it’s easy to see that the cross section curve $\gamma$ at one of the necks where surgery was performed on $S_{T}^{-}$ (the presurgery surface) is homotopically nontrivial in $M_{T}$. The surgery shows that $\gamma$ bounds a disc in $N$ though, so that $S_{T}^{-}$ is not algebraically incompressible. Since $T$ was the least such time, this implies that $M$ itself is not algebraically incompressible, which gives a contradiction.